OpenStudy (anonymous):

Alicia estimated her test score to be 85. Her actual score was 80. What was the percent error in Alicia’s estimate? Round your answer to the nearest tenth of a percent.

OpenStudy (anonymous):

@Michele_Laino @MissNoName @mathmath333 @Mehek14 @Abhisar @AlexandervonHumboldt2

OpenStudy (anonymous):

6

OpenStudy (anonymous):

6%

OpenStudy (anonymous):

5.9%

OpenStudy (michele_laino):

the requested error, is given by the subsequent computation: \[\huge \frac{{85 - 80}}{{85}} \times 100 = ...?\]

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \frac{ \rm |correct-estimated| }{\rm correct}\times 100 }\)

OpenStudy (solomonzelman):

Isn't that the case?

OpenStudy (anonymous):

5.8

OpenStudy (solomonzelman):

We are searching for how much wrong was the estimated score from the actual score... (not the other way around) I might be wrong, but I think we are dividing by 80.

OpenStudy (michele_laino):

It is \(5.88 \%\) so, that's right!

OpenStudy (anonymous):

k

OpenStudy (michele_laino):

which can be rounded to \(6 \%\)

OpenStudy (anonymous):

Round your answer to the nearest tenth of a percent.

OpenStudy (anonymous):

5.7

OpenStudy (michele_laino):

sorry, it can be rounded to \(5.9 \%\)

OpenStudy (anonymous):

yeah i made a mistake

OpenStudy (solomonzelman):

are you sure we are dividing by 85, not 80?

OpenStudy (solomonzelman):

I'm not very certain about this, that's why I'm asking.